OpenStudy (anonymous):

I GIVE MEDALS. The soccer player is running at a speed of 4 m/s when he slides on the ground, coming to a stop in a distance of 2 m. What was his acceleration?

OpenStudy (aravindg):

u=4 v=0 S=2 Use v^2-u^2= 2aS

OpenStudy (anonymous):

so it's 0^2-4^2=2a(2)? how would i solve for acceleration? @AravindG

OpenStudy (aravindg):

Simplify on both sides

OpenStudy (anonymous):

do i need tom move the equation around to make it where it says "a="? @AravindG

OpenStudy (aravindg):

Yep

OpenStudy (aravindg):

You could divide by 4 on both sides to get a

OpenStudy (anonymous):

okay, so i got this far. 0^2-4^2=2a(2) =-16=2a(2) =-4=2a(2) is this correct so far? how would i solve for the other side? @AravindG

OpenStudy (aravindg):

Yeah right Now divide by 4 on both sides

OpenStudy (anonymous):

how would divide the right side with a variable in it? @AravindG

OpenStudy (aravindg):

\[\dfrac{-4}{4}=\dfrac{4a}{4}\]

OpenStudy (aravindg):

Simple as that

OpenStudy (anonymous):

so the final answer would be -1=1a? @AravindG

OpenStudy (aravindg):

Yep ie a=-1 :)

OpenStudy (anonymous):

thanks so much! @AravindG

OpenStudy (aravindg):

yw :)